OpenStudy (i_love_my_nieces):

Find the range of the function f(x)=4x-1 for the domain {-1, 0, 1, 2, 3}.

OpenStudy (i_love_my_nieces):

@Michele_Laino

OpenStudy (michele_laino):

the requested range, is given by the subsequent set: \[\Large range = \left\{ {f\left( { - 1} \right),f\left( 0 \right),f\left( 1 \right),f\left( 2 \right),f\left( 3 \right)} \right\}\]

OpenStudy (michele_laino):

for example, we can write this: \[\Large f\left( 0 \right) = \left( {4 \cdot 0} \right) - 1 = - 1\]

OpenStudy (i_love_my_nieces):

Carry on

OpenStudy (i_love_my_nieces):

What is the answer

OpenStudy (michele_laino):

please compute these quantities: \[\Large \begin{gathered} f\left( { - 1} \right) = \left( {4 \cdot \left( { - 1} \right)} \right) - 1 = ...? \hfill \\ \hfill \\ f\left( 1 \right) = \left( {4 \cdot 1} \right) - 1 = ...? \hfill \\ \hfill \\ f\left( 2 \right) = \left( {4 \cdot 2} \right) - 1 = ...? \hfill \\ \hfill \\ f\left( 3 \right) = \left( {4 \cdot 3} \right) - 1 = ...? \hfill \\ \end{gathered} \]

OpenStudy (i_love_my_nieces):

idk how to I need the answer I have to get out of where I am so can you tell me please

OpenStudy (anonymous):

Answer : Range is (-5,-1,3,7,11)

OpenStudy (i_love_my_nieces):

can you help with some more?